If demand for pork is given by: QD = 200 – 6P + 2Y, when the...

  1. If demand for pork is given by: QD = 200 – 6P + 2Y, when the price of pork is £8, a rise inconsumers’ income from £100 to £150 leads to:

    1. a) a fall in demand and an income elasticity of -0.14, pork is an inferior good

    2. b) a rise in demand and an income elasticity of 0.14, pork is a normal good and a necessity

    3. c) a rise in demand and an income elasticity of 7.08, pork is a luxury good

    4. d) a fall in demand and an income elasticity of -7.08, pork is an inferior good.

  2. what is 2Y and it doesnt say what is the new QD? WHAT DO WE DO
Answer #1

Income elasticity of Demand: is defined a change in quantity demanded due to change in income.

Formula - Ei= % change in Q/ % change in Income = (New Qd - Old Qd / Old Qd )/ ((New Y- Old Y)/ Old Y)

Ei= 0-1 normal good

Ei<0 - inferior good

Ei>1 - luxury good

According to the question,

QD = 200 – 6P + 2Y

Demand function is a decreasing function in prices and increasing function in income.

Price = £8

Old Income - £100

New Income - £150

Old Demand = 352 units

New Demand = 452 units

Income Elasticity = Ei= % change in Q/ % change in Income = (452-352)/352/((150-100)/100)

= (100/352)/ (50/100)

=0.56

So, the demand increases and it is a normal good with income elasticity is 0.56.

Change in Quantity is because of icnrease in income

LDOeAG1K69kAAAAASUVORK5CYII=